OpenStudy (anonymous):

A water ride at a local water park has a ride shaped like a cone that acts like a funnel whereby guests swirl around the cone until they drop through its center. There is one ride for adults and a similar, smaller version for children. If the adult ride has a radius of 32 feet and the child ride has a radius of 24 feet, what is the ratio between the volumes of each ride?..

OpenStudy (anonymous):

I'M assuming the height (h)is the same volume = 1/3 pi r^2 h so adults = 1/3 pi 32^2 h 1024/3 pi h kids =1/3 pi 24^2 h =576/3 pi h RATIO kids to adults 576/3 pi h = 1024/3 pi h cancel the h on both side 576/3 pi = 1024/3 pi cancel pi on both sides 576/3 =1024/3 multiple by 3 576=1024 work the ratio down dividing by 2's and 3's ect 36:64 18:32 9:16